OpenStudy (anonymous):

true or false this set of three numbers the measurements of a right triangle 25,23,7

OpenStudy (phi):

if it is a right triangle then the pythagorean theorem applies: \[ a^2 + b^2 = c^2\] where a and b are the legs and c is the hypotenuse (longest side) 25 is your longest side, so we should test to see if \[ 7^2 + 23^2 =? \ 25^2 \] if this is true, you have a right triangle... if not true, then you don't have a right triangle.